OpenStudy (anonymous):

Can someone please tell me if I'm correct? f(x) = 3x^2 + x^3 -2 and g(x) -2 + 3x^2 + x^2 are NOT equal?

OpenStudy (anonymous):

I think they are equal

OpenStudy (anonymous):

You would replace x and y variable so it would be "x=3y^2+x^3+x^2, 3y^2+y^3 = 10 and 10-2 = 8

OpenStudy (anonymous):

for the second problem:

OpenStudy (anonymous):

x=-2+3y^2+y^2 and 3y^2 + 1^2 = 10 - 2 = 8

OpenStudy (anonymous):

Makes sense to me. Thank you